OpenStudy (anonymous):

Find the distance between the points (-1, 3) and (5, 0).

OpenStudy (anonymous):

\[d = \sqrt{(-1-5)^2 + (3-0)^2}\]

OpenStudy (anonymous):

wi

OpenStudy (anonymous):

\[\sqrt{45}?\]

OpenStudy (anonymous):

\[\sqrt{(y2-y1)^2+(x2-x1)}\] This is the formula we use to find the distance between points. So if we pluck (-1,3) & (5,0) in this formula we have \[\sqrt{(-1-5)^2+(3-0)^2}\] Then we do what are in bose parenthesis and get- \[\sqrt{(-6)^2+(3)^2}\] Once we put the exponents & add it together we now have- \[\sqrt{45}\]

OpenStudy (anonymous):

There's not a choice in that

OpenStudy (anonymous):

if you round it you get 7 if that helps

OpenStudy (anonymous):

 Those are the choices